OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4,-2) and perpendicular to y=-x+6

OpenStudy (larseighner):

The slope of a perpendicular line to a line with slope m is -1/m. Plug into point-slope formula with the point you have.

OpenStudy (anonymous):

the genral eq of a point slope form is (y-y1)=m(x-x1) where m= slope (x1,y1)=(-4,-2) and if the line is perp. to that one..then product of thier slopes must be equal -1. use this to find slope=m=-y/x

OpenStudy (anonymous):

Ok so Y-(-2)=-1(x--4) Then do I turn -2 and -4 into positive ?

OpenStudy (anonymous):

yes you can...but 1st find the slope...

OpenStudy (anonymous):

since the equation is perpendicular to y=-x+6, slope of this equation is -1. since for perpendicular line, m1m2=-1, thus the slope of new equation becomes 1. since it passes (-4,-2) y+2=1(x+4) y+2=x+4 x-y+2=0 thats your equation (y) dont forget to give the medal (y)

OpenStudy (larseighner):

The slope of the given line is -1. So the slope of the perpendicular is - 1/(-1) = 1